# On the second Bohr radius for vector valued pluriharmonic functions

### v1

@@ 3.2. Second Bohr radii for homogeneous plynomials
In light of Theorem 3.1, we deduce the following asymptotic estimates for Banach sequence spaces, which have wide applications.
@@
    $L_{\lambda}(B_{Z_{n}},p,X)\geq\begin{cases}E_{1}(X)\,\frac{1}{s\,t\,n^{\frac{1}{q}}}\,\max\left\{\frac{1}{\sqrt{n}\,\left\lVert Id:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert},\,\frac{1}{e\,\left\lVert\sum_{k=1}^{n}e^{*}_{k}\right\rVert_{Z^{*}}}\right\}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}&\text{for $p=1$},\\[5.69054pt] E_{3}\,\frac{1}{s}\,\left\lVert I:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert^{-\frac{2}{p}}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,&\text{for $p\geq 2$},\\[5.69054pt] \frac{E_{2}(X)}{s\left(tn^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}}\,\max\left\{\frac{1}{(\sqrt{n})^{1-\theta}\left\lVert Id:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert},\frac{\left\lVert I:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert^{-\theta}}{e^{(1-\theta)}\left\lVert\sum_{k=1}^{n}e^{*}_{k}\right\rVert_{Z^{*}}^{1-\theta}}\right\}\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\hskip-8.53581pt&\text{for $1<p<2$}.\end{cases}$
@@
We start by recalling two important facts from the theory of finite-dimensional Banach spaces with unconditional bases. Let $W=(\mathbb{C}^{n},||.||)$ be a Banach space whose canonical unit vectors $e_{k}$ form a normalized $1$-unconditional basis. Then $\left\lVert Id:W\rightarrow\ell_{2}^{n}\right\rVert=\sup_{z\in B_{W}}(\sum_{m=1}^{n}|z_{m}|^{2})^{1/2}$. Moreover, $\left\lVert Id:W\rightarrow\ell_{1}^{n}\right\rVert=\sup_{z\in B_{W}}\sum_{m=1}^{n}|z_{m}|=\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}$ for all $z\in W$.
@@
We now turn to the proof of part $(2)$. By [32, Theorem 1.d.5], we may assume without loss of generality that $M^{(2)}(X)=1$, which in turn implies $\left\lVert Id:\ell_{2}^{n}\rightarrow Z_{n}\right\rVert=1$. We next recall an important estimate from [19, (5.3), p. 191]: if $W=(\mathbb{C}^{n},||.||)$ is a symmetric Banach space satisfying $M^{(2)}(W)=1$, then
(3.13) $\left\lVert Id:W\rightarrow\ell^{n}_{2}\right\rVert=\frac{\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}}{\sqrt{n}}.$

### v2

@@ 5. Applications to Banach sequence spaces
As an application of Theorem 3.1, we deduce the following asymptotic estimates for Banach sequence spaces in finite dimensional settings.
@@
    $L_{\lambda}(B_{Z_{n}},p,X)\geq\begin{cases}E_{1}(X)\,\frac{1}{s\,t\,n^{\frac{1}{q}}}\,\max\left\{\frac{1}{\sqrt{n}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert},\,\frac{1}{e\,\left\lVert\sum_{k=1}^{n}e^{*}_{k}\right\rVert_{Z^{*}}}\right\}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}&\text{for $p=1$},\\[5.69054pt] E_{3}\,\frac{1}{s}\,\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert^{-\frac{2}{p}}\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,&\text{for $p\geq 2$},\\[5.69054pt] \frac{E_{2}(X)}{s\left(tn^{\frac{1}{q}}\right)^{\frac{(2-p)}{p}}}\,\max\left\{\frac{1}{(\sqrt{n})^{1-\theta}\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert},\frac{\left\lVert\mathrm{Id}:\ell^{n}_{2}\rightarrow Z_{n}\right\rVert^{-\theta}}{e^{(1-\theta)}\left\lVert\sum_{k=1}^{n}e^{*}_{k}\right\rVert_{Z^{*}}^{1-\theta}}\right\}\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\hskip-8.53581pt&\text{for $1<p<2$}.\end{cases}$
@@
We start by recalling two important facts from the theory of finite-dimensional Banach spaces with unconditional bases. Let $W=(\mathbb{C}^{n},||.||)$ be a Banach space whose canonical unit vectors $e_{k}$ form a normalized $1$-unconditional basis. Then $\left\lVert\mathrm{Id}:W\rightarrow\ell_{2}^{n}\right\rVert=\sup_{z\in B_{W}}(\sum_{m=1}^{n}|z_{m}|^{2})^{1/2}$. Moreover, $\left\lVert\mathrm{Id}:W\rightarrow\ell_{1}^{n}\right\rVert=\sup_{z\in B_{W}}\sum_{m=1}^{n}|z_{m}|=\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}$ for all $z\in W$.
@@
We now turn to the proof of part $(2)$. By [32, Theorem 1.d.5], we may assume without loss of generality that $M^{(2)}(X)=1$, which in turn implies $\left\lVert\mathrm{Id}:\ell_{2}^{n}\rightarrow Z_{n}\right\rVert=1$. We next recall an important estimate from [19, (5.3), p. 191]: if $W=(\mathbb{C}^{n},||.||)$ is a symmetric Banach space satisfying $M^{(2)}(W)=1$, then
(5.1) $\left\lVert\mathrm{Id}:W\rightarrow\ell^{n}_{2}\right\rVert=\frac{\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}}{\sqrt{n}}.$